OpenStudy (anonymous):

Solve for x and y A. 3x+8y=2 B. x=2n -3x-2y=-5 y= 5n 2x+y=54

jhonyy9 (jhonyy9):

3x+8y=2 -3x-2y=-5 assume ------------ 0 6y=-3 y= -3/6 = -1/2 3x + 8(-1/2) =2 3x -4 =2 3x=6 x=6/3 =2 x=2 y= -1/2

jhonyy9 (jhonyy9):

B 2(2n) +5n=54 4n+5n=54 9n=54 n=54/9=6 n=6 x=2n=12 y=5n=30